OpenStudy (anonymous):

find the distance between (0,9) and (-40,0). express answer in simplest form.

OpenStudy (anonymous):

\[\sqrt{40^2+9^2}\] \[\sqrt{1600+81}\] \[\sqrt{1681}\]

OpenStudy (anonymous):

so 41?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank u